OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3? answers x - y + 2 = 0 x + y + 2 = 0 x + y - 2 = 0

OpenStudy (lgbasallote):

it is parallel to y = -x+3 so slope of parallel lineis -1 use point slope form y - (-1) = -1(x - (-1)) y + 1 = -(x+!) y + 1 = -x - 1 y = -x - 1 - 1 y = -x - 2

OpenStudy (lgbasallote):

or rewrite as x +y + 2 = 0

OpenStudy (anonymous):

thank you